OpenStudy (anonymous):

Sara is wrapping gift boxes. Each gift box is a rectangular prism with a square base of side 6 inches. The difference between the heights of the two gift boxes is 3 inches. If the larger box has a surface area of 216 square inches, what is the difference in the surface area of the two boxes? 108 square inches 72 square inches 36 square inches 18 square inches

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@dpaInc

OpenStudy (anonymous):

i think im doing something wrong because i got 144

jimthompson5910 (jim_thompson5910):

The base is 6^2 = 36 square inches So the area of the two bases combined is 2*36 = 72 square inches

jimthompson5910 (jim_thompson5910):

The larger box has a surface area of 216 sq inches, so it has a lateral surface area of 216 - 72 = 144 square inches

jimthompson5910 (jim_thompson5910):

You can then use this lateral surface area to find the height of the larger box

OpenStudy (anonymous):

6?

OpenStudy (anonymous):

therefore it is...72?

jimthompson5910 (jim_thompson5910):

good, so the smaller box has a height of 3 inches (since the larger box is taller by 3 inches)

OpenStudy (anonymous):

yeah

jimthompson5910 (jim_thompson5910):

the lateral surface area of the smaller box is then....

OpenStudy (anonymous):

72 right? because 24*3

jimthompson5910 (jim_thompson5910):

yep, so the total surface area of the smaller box is 72 + 72 = 144 you add the base area (72) and the lateral surface area (72) to get 144

jimthompson5910 (jim_thompson5910):

the difference of the two surface areas is then Big - Small = 216 - 144 = 72

OpenStudy (anonymous):

ohhh i actually didnt read the question properly

OpenStudy (anonymous):

i thought the question was asking what the area of the smaller one was

OpenStudy (anonymous):

ok thanks

jimthompson5910 (jim_thompson5910):

alternatively, you can draw out the two boxes, then just focus on the upper half of the larger box that has a height of 3 and if you just focus on the lateral surface area and unfold it out, you'd get this 

jimthompson5910 (jim_thompson5910):

and you'd get the area to be 3*24 = 72 square inches

OpenStudy (anonymous):

ok thx

OpenStudy (anonymous):

so answer is B right? just double checking..

jimthompson5910 (jim_thompson5910):

yep